# Turing’s Test, a Beautiful Thought Experiment

### v2

@@ Abstract.
In the wake of large language models, there has been a resurgence of claims and questions about the Turing test and its value for AI, which are reminiscent of decades of practical “Turing” tests. If AI were quantum physics, by now several “Schrödinger’s” cats could have been killed. Better late than never, it is time for a historical reconstruction of Turing’s beautiful thought experiment. In this paper I present a wealth of evidence, including new archival sources, give original answers to several open questions about Turing’s 1950 paper, and address the core question of the value of Turing’s test.
Alan Turing, Turing test, Thought experiment, Foundations of AI & computer science, Galileo Galilei, History of science, History of AI
journalyear: 2024doi: XXXXXXX.XXXXXXX
Figure 1. Alan Turing (1912-1954). Photographs of Alan Turing, copyright The Provost and Scholars of King’s College Cambridge 2023. Archives Centre, King’s College, Cambridge, AMT/K/7/12. Reproduced with permission.
1. Introduction
@@
In recent decades, the Turing test has been used as a practical experiment for publicity purposes and aptly criticized (Shieber, 1994a, b; Vardi, 2014), and has been the whipping boy of AI (Hayes and Ford, 1995), cognitive sciences and analytic philosophy (Shieber, 2004), and increasingly, with the rise of AI, the humanities and social sciences (Brynjolfsson, 2022). It is not uncommon for criticism coming from all these areas to take Turing’s test literally, assuming that he encouraged deception as a criterion and/or proposed a crucial experiment to establish the existence of machine intelligence.
Now, in the wake of large language models, science and technology outlets ask whether Turing’s test can be a ‘benchmark’ for AI (Biever, 2023), and whether it is ‘dead’ (Wells, 2023). Based on recent primary research (Gonçalves, 2023e, a, b, c, d), in this paper I present a mass of evidence, including newly discovered archival sources, and a new perspective on Turing’s test.
I address a few problems that we will keep track as we go along: ($P_{1}$) Why would Turing design a test based on imitation, which can be seen as encouraging deception? ($P_{2}$) Why to present multiple versions of a test, as opposed to a well-defined, controlled experiment? ($P_{3}$) Why gender imitation in a test of machine intelligence, and why conversation for an intelligence task? ($P_{4}$) Why are practical “Turing” tests circular, as Hayes and Ford (Hayes and Ford, 1995) claimed? Finally, the core problem, ($P_{5}$) What is the value of Turing’s test for AI?
Section 2 presents a reading of Turing’s 1950 paper. Section 3 introduces newly discovered archival sources and examines Turing’s concept of imitation, addressing $P_{1}$. Section 4 shows that his presentation of the imitation game fits “the basic method of thought experiments” (Mach, 1897), addressing $P_{2}$. Section 5 reconstructs the historical conditions of Turing’s proposal, addressing $P_{3}$. Section 6 draws a parallel with the history of “the most beautiful experiment in the history of science” (Palmieri, 2005), addressing $P_{4}$. Finally, Section 7 revisits the history of AI, addressing $P_{5}$, and Section 8 concludes the paper.
2. What is the Turing Test?
@@
In 1950, Alan Turing (Fig. 1) published the second of his three seminal papers, ‘Computing machinery and intelligence’ (Turing, 1950). The text has 28 pages, divided in seven sections, §1-§7. Three main logical steps can be identified in his argument: the proposal (§1-§3, 3+ pp.), the science (§4-§5, 6 pp.), and the discussion (§6-§7, 18+ pp.).
@@
The proposal sought to replace the question “Can machines think?,” which he considered “too meaningless to deserve discussion” (p. 442), [^1] Turing was coming from unstructured multidisciplinary debates in at least two editions of a seminar, “Mind and Machine,” held at the Philosophy Department of Manchester University in October and December, 1949. Of the latter, one participant wrote in a Christmas postcard sent to Warren McCulloch: “I wish you had been with us a few days ago we had an amusing evening discussion with Thuring [sic], Williams, Max Newman, Polyani [sic], Jefferson, J Z Young & myself $\ldots$ An electronic analyser and a digital computer (universal type) might have sorted the arguments out a bit.” Jules Y. Bogue to McCulloch, c. December 1949. American Philosophical Society, Warren S. McCulloch Papers, Mss.B.M139_005. Thanks to J. Swinton for this archival finding. with the imitation game. The purpose was to change the common meaning of the word ‘machine’ (e.g., a steam engine, a bulldozer) in light of the new mathematical science of ‘universal’ digital computing. The imitation game would allow for a grounded discussion of ‘machine’ and ‘thinking,’ seeking to expand the meaning of ‘thinking’ and detach it from the human species, much as the meaning of ‘universe’ was once detached from the Earth.
@@
In 1950, one of the OED definitions of ‘machine’ was: [^2] New English Dictionary. Oxford, Vol. VI, Part II, M-N, p. 7. “a combination of parts moving mechanically as contrasted with a being having life, consciousness and will $\ldots$ Hence applied to a person who acts merely from habit or obedience to a rule, without intelligence, or to one whose actions have the undeviating precision and uniformity of a machine.” Thus, by definition, common sense did not allow the meanings of ‘machine’ and ‘thinking’ to overlap. Despite Turing’s emphasis in his opening paragraph that he did not intend to discuss how these words were “commonly used” (p. 433), the hostility to his proposal can be seen from one of the first reactions, from a participant in the 1949 Manchester seminars, who quoted the above OED definition to appeal to common sense (Mays, 1952).
@@
The new question, which Turing considered to have a “more accurate form” (Turing, 1950, p. 442), would be based on a vivid image, his “criterion for ‘thinking”’ (p. 436), which he called interchangeably the ‘imitation game’ and his ‘test.’ [^3] For Turing’s exact references to his test in all known sources, see (Gonçalves, 2023e, p. 2). The new question is whether a machine, playing A, the deceiver, can imitate a woman, a man, a human being, or another machine, playing B, the assistant, in a remotely played conversation game, and deceive an average interrogator, playing C, the judge, about its machine condition.
However, the details and exact conditions of the imitation game as an experiment slipped through Turing’s text in a series of variations that defies interpretation. A structural reading of the text identifies four different conditions of the game with respect to players A-B, namely, man-woman (p. 433), machine-woman (p. 434), machine-machine (p. 441), and machine-man (p. 442). These different conditions relate to four variants of the “new” question that Turing posed to replace his “original” question (see Box 1). In addition to varying the genus/species (types) of the players, he also increased the storage and speed of the machine and provided it with a hypothetically appropriate program ($Q^{\prime\prime\prime}$), and suggested a base time for the interrogation session ($Q^{\prime\prime\prime\prime}$). Other seemingly relevant parameters were not mentioned, such as the number of interrogators used to arrive at a statistically sound conclusion, although their profile is mentioned — they should be “average” —, and later reiterated — they “should not be expert about machines.”444‘Can automatic calculating machines be said to think?’, Broadcast on BBC Third Programme, 14 and 23 Jan. 1952. Archives Centre, King’s College, Cambridge, AMT/B/6.
BOX 1
Turing’s various questions and conditions
@@
$Q^{\prime}$: “What will happen when a machine takes the part of A in this [man-woman] game?’ Will the interrogator decide wrongly as often when the game is played like this [machine-woman] as he does when the game is played between a man and a woman? These questions replace our original, ‘Can machines think?’ ” (pp. 433-434) $Q^{\prime\prime}$: “There are already a number of digital computers in working order, and it may be asked, ‘Why not try the experiment straight away? It would be easy to satisfy the conditions of the game. A number of interrogators could be used, and statistics compiled to show how often the right identification was given.’ The short answer is that we are not asking whether all digital computers would do well in the game nor whether the computers at present available would do well, but whether there are imaginable computers which would do well.” (p. 436) $Q^{\prime\prime\prime}$: “It was suggested tentatively that the question [$Q$] should be replaced by [$Q^{\prime\prime}$] $\ldots$ But in view of the universality property we see that either of these questions is equivalent to this, ‘Let us fix our attention on one particular digital computer C. Is it true that by modifying this computer to have an adequate storage, suitably increasing its speed of action, and providing it with an appropriate programme, C can be made to play satisfactorily the part of A in the imitation game, the part of B being taken by a man?’ ” (p. 442) $Q^{\prime\prime\prime\prime}$: “I believe that in about fifty years’ time it will be possible to programme computers, with a storage capacity of about $10^{9}$, to make them play the imitation game so well that an average interrogator will not have more than 70 per cent. chance of making the right identification after five minutes of questioning.” (p. 442)
3. Imitation: from 1936 to 1950
Because the machine must imitate stereotypes of what it is not, Turing’s proposal has often been criticized for encouraging fakes and tricks. But this is part of the literal reading that Turing would have meant his test as a practical experiment about short conversations, gendered machines, and how to fine-tune them to fool average human interrogators. It misses the point of Turing’s irony (Gonçalves, 2023b) and that his notion of imitation in 1950 was largely in continuity with his 1936 paper (Turing, 1936), as hinted at in his 1947 lecture,555‘Lecture to L.M.S. Feb. 20 1947.’ Archive Centre, King’s College, Cambridge, AMT/B/1. and as newly discovered correspondence with the Mexican-Canadian computer pioneer Beatrice Worsley (1921-1972) helps to clarify (see Box 2).
In his letter to Worsley, Turing seems to be more interested in the relations between “the motions” of Turing machines and infinite computers, whose behavior can be non-periodic. Perhaps he thought of the living human brain as an infinite computer, in the sense that it has a continuous interface with its environment, which constantly intervenes and changes its logical structure. [^6] Cf. “Intelligent Machinery”, written in 1948 as a technical report to the National Physical Laboratory. Archives Centre, King’s College, Cambridge, AMT/C/11. Now, the imitation game puts into empirical form the relation between digital computers, whose behavior is ultimately periodic, and the behavior of the human players. Can the behavior of their brains be approximated by a digital computer? Turing pursued this question. For his May 1951 broadcast, he wrote: “the view which I hold myself, that it is not altogether unreasonable to describe digital computers as brains $\ldots$ If it is accepted that real brains, as found in animals, and in particular in men, are a sort of machine it will follow that our digital computer, suitably programmed, will behave like a brain.”777‘Can digital computers think?’, broadcast on BBC Third Programme, 15 May 1951. Archives Centre, King’s College, Cambridge, AMT/B/5.
Even if the human brain can only be compared to an infinite computer, could it not be simulated by a digital computer equipped with a sufficiently large memory? An excerpt of another newly discovered Turing letter to Worsley from mid-1951 can give more contour and provide further insight into Turing’s views (see Box 3).
BOX 2
@@ 3. Imitation: from 1936 to 1950
Yours sincerely, A. M. Turing888 Turing to B. H. Worsley, June 11, 1951, typeset; emphasis added. Unpublished writings of Alan Turing, copyright The Provost and Scholars of King’s College Cambridge 2023. B.H. Worsley Collection, Archives Center, National Museum of American History, Smithsonian Institution. Quoted with permission. Thanks to Mark Priestley for finding and kindly sharing this source.
A highlight in this excerpt is Turing’s view that to the extent that the behavior of a neuron can be described as a stochastic process, it would be possible to “calculate random samples” of the mechanism that embodies the brain and then imitate it. [^9] Note the connection with the imitation game and his wartime experience studying and imitating the behavior of the Enigma machines used by the Nazi forces (Hodges, 1983). An effective imitation of the brain by a machine would require knowledge of the anatomy and physiology of the brain to inspire an appropriate program, as well as much more storage and speed than was available to the Ferranti Mark I at the time (see Fig. 2). Another important element in the excerpt is Turing’s point that, even if a thinking machine is possible, the relation he has in mind is not one of identity but one of analogy: “It’ll just be another species of the thinking genus.”
An original answer to problem $P_{1}$, why design a test based on imitation, which can be seen as encouraging deception, is as follows. Actually, imitation was Turing’s fundamental principle of the new science of universal digital computing. He conceived his 1950 paper largely in continuity with his 1936 paper. Both were based on his core concepts of machine and imitation, i.e., what it takes for a machine to imitate another machine. A point of difference is that by 1950 he had generalized the machine architecture and how ‘universal’ imitation can be achieved. Using Turing’s 1948 language,101010“Intelligent Machinery,” op. cit. universality could be achieved with an ‘organized’ machine (1936), or with an ‘unorganized’/‘learning’ machine (1948/1950). Whereas in 1936 the machine would be given an a priori, specified, and fixed table of instructions for each task, in 1950 it would also be able to perform a new task by changing its logical structure as a result of learning from experience, much as the brain does, “by changing its neuron circuits by the growth of axons and dendrites.” [^11] Turing to Ross Ashby, circa November 19, 1946. British Library, Collection ‘W. Ross Ashby: correspondence of W. Ross Ashby’, Add MS 89153/26.
The kind of forgery and trickery that occurs in publicity-oriented practical “Turing” tests has nothing to do with Turing’s 1950 proposal. In 1951, he warned: “It would be quite easy to arrange the [machine’s] experiences in such a way that they automatically caused the structure of the machine to build up into a previously intended form, and this would obviously be a gross form of cheating, almost on a par with having a man inside the machine.”121212‘Intelligent machinery, a heretical theory’, a lecture given to ‘51 Society’ at Manchester, c. 1951. Archives Centre, King’s College, Cambridge, AMT/B/4. The “human fallibility” that Turing encouraged the machine to show was not meant to be artificially introduced, but rather a by-product of learning from experience. This 1950 passage clarifies: “Another important result of preparing our machine for its part in the imitation game by a process of teaching and learning is that ‘human fallibility’ is likely to be omitted in a rather natural way, i.e., without special ‘coaching’ ” (Turing, 1950, p. 459). That is, for the machine to be a valid player, it must not be specially prepared for the test. This is one of the reasons why we have never seen a practical Turing test.
BOX 3
Simulation of neuron processes
@@
I do not think you will be able to find any clue to essential differences between brains and computing machines (if there are any), in neuron behaviour. So long as what we know about a neuron can be embodied in the description of stochastic processes, the behaviour of any mechanism embodying such neurons can, in principle, be calculated by a suitable enlarged and speeded up Ferranti [Mark II] machine.131313‘Ferranti’ is typed and erased, and ‘Mark II’ is added in pencil, referring to a version of the evolving Manchester electronic computer. More accurately I should say that one can calculate random samples of its behaviour. I think any attempt to draw any sharp line between what machine and brain can do will fail. I think it is largely a quantitative matter. Probably one needs immensely more storage capacity then [sic] we have got, and possibly more than we shall ever have. Perhaps we may have enough capacity, but just won’t find an appropriate programme. Naturally one won’t make a man that way ever. It’ll just be another species of the thinking genus.
Yours sincerely, A. M. Turing141414 Turing to B. H. Worsley, circa June, 1951, Turing’s emphasis. Credit for this source is exactly the same as for that of Box 2. Quoted with permission.
Figure 2. Control console of Ferranti Mark I and a group with Turing’s secretary at the Computing Machine Laboratory, Sylvia Robinson, pretending to play chess with the machine, c. 1955. Courtesy of The University of Manchester.
4. The method of thought experiments
The various rhetorical questions Turing posed, $Q^{\prime}\ldots Q^{\prime\prime\prime\prime}$, to replace the original question, $Q$, can be generalized as follows (Gonçalves, 2023e):
@@ 4. The method of thought experiments
It has been largely unnoticed that the various questions instantiating $Q^{\star}$ follow a case-control methodology, applied in two stages. At the more obvious intra-game level, A plays the case, and B plays the control. However, at the inter-game level, two variants set the case (machine-woman and the machine-man) and the other two set the control (man-woman and the machine-machine). While the first two are open, creating suspense around the test, the latter two are resolved as follows. It is known that a man (A) can possibly imitate gender stereotypes associated with a woman (B) to the point of deceiving an interrogator (C) despite their physical differences. This is the very premise of the parlor games that existed at the time. Further, regarding the machine-machine variant, it is also known that a digital computer (A), because of its universality property, as Turing explained in the paper (Turing, 1950, §§4, 5), can possibly imitate any discrete-state machine (B), despite their architectural differences.
We can now explore how Turing’s presentation of his test conforms to Ernst Mach’s reconstruction of “the basic method of thought experiments,” which is variation, continuously if possible. Mach is the author of perhaps the most classic text on thought experiments in the modern scientific tradition (Mach, 1897), in which he developed observations and insights based on a wealth of examples from the history of modern physics, mathematics, and common sense experience. He wrote: “By varying the conditions (continuously if possible), the scope of ideas (expectations) tied to them is extended: by modifying and specializing the conditions we modify and specialize the ideas, making them more determinate, and the two processes alternate” (p. 139). Mach illustrated his point with the process of discovery of universal gravitation (Mach, 1897, pp. 138-139):
A stone falls to the ground. Increase the stone’s distance from the earth, and it would go against the grain to expect that this continuous increase would lead to some discontinuity. Even at lunar distance the stone will not suddenly lose its tendency to fall. Moreover, big stones fall like small ones: the moon tends to fall to the earth. Our ideas would lose the requisite determination if one body were attracted to the other but not the reverse, thus the attraction is mutual and remains so with unequal bodies, for the cases merge into one another continuously $\ldots$ discontinuities are quite conceivable, but it is highly improbable that their existence would not have betrayed itself by some experience. Besides, we prefer the point of view that causes less mental exertion, so long as it is compatible with experience.
@@
Turing’s imitation game extended the scope of ideas and expectations established earlier in his 1936 paper, moving from machine-machine and restricted human-machine imitation in 1936151515“We may compare a man in the process of computing a real number to a machine which is only capable of a finite number of conditions” (Turing, 1936, p. 231). to more general human-machine imitation in 1950.

### v3

# turing-argument.bib

@@ Abstract
In the wake of the latest trends of artificial intelligence (AI), there has been a resurgence of claims and questions about the Turing test and its value, which are reminiscent of decades of practical “Turing” tests. If AI were quantum physics, by now several “Schrödinger’s” cats would have been killed. It is time for a historical reconstruction of Turing’s beautiful thought experiment. This paper presents a wealth of evidence, including new archival sources, and gives original answers to several open questions about Turing’s 1950 paper, including its relation with early AI.
@@
Since the early 1990s, Turing’s test has been used for publicity purposes as a practical experiment, and has been aptly criticized [shieber1994, shieber1994b, vardi2014]. It has been the whipping boy of AI [hayes1995], cognitive sciences and analytic philosophy (cf. [shieber2004]), and increasingly, with the rise of AI, the humanities and social sciences [brynjolfsson2022]. It is not uncommon for critics from all these fields to take Turing’s test literally, while ignoring parts of his 1950 text. It is often assumed that he was promoting deception as a criterion for intelligence and/or proposing a crucial experiment to establish the existence of machine intelligence. Now, with the latest trends of AI technology, publications ask whether the Turing test can be a “benchmark” for AI [biever2023], and whether it is “dead” [wells2023]. Based on recent primary research [goncalves2023experiment, goncalves2023galileo, goncalves2023irony, goncalves2023lovelace, goncalves2023argument], this paper presents a new perspective on Turing’s test.
Figure 1: Alan Turing (1912-1954). Photographs of Alan Turing, copyright The Provost and Scholars of King’s College Cambridge 2023. Archives Centre, King’s College, Cambridge, AMT/K/7/12. Reproduced with permission.
A structural reading of Turing’s 1950 paper is presented. Newly discovered archival sources are introduced and Turing’s concept of imitation is examined. It is argued that his discursive presentation of the imitation game in various versions, as opposed to a well-defined, controlled experiment, conforms to what the physicist and historian Ernst Mach called “the basic method of thought experiments” [mach1897]. The historical conditions of Turing’s proposal are reconstructed from a controversy about mind and computing machine in England, contextualizing his requirement of gender imitation and conversational performance as success criteria for machine intelligence. Turing’s responses to critics and propaganda for machine intelligence are highlighted. Finally, the question of the value of Turing’s test and its relation to the history of AI is revisited.
WHAT IS THE TURING TEST?
@@
In 1950, Alan Turing (Fig. 1) published “Computing Machinery and Intelligence” [turing1950], the second of his three seminal papers. [^1] The other two being “On Computable Numbers” [turing1936] and “The Chemical Basis of Morphogenesis” [turing1952b]. The text has 28 pages, divided in seven sections, §1-§7. Three main logical steps can be identified in his argument: the proposal (§1-§3, in $3+$ pp.), the science (§4-§5, in 6 pp.), and the discussion (§6-§7, in 18+ pp.). This structure, with its thematic order and size distribution, can be revealing about the nature of Turing’s paper and argument.
@@
The proposal sought to replace with the imitation game the question “Can machines think?,” which he considered “too meaningless to deserve discussion” (p. 442; Turing was coming from unstructured multidisciplinary debates in two editions of a seminar, “Mind and Machine,” held at the Philosophy Department of Manchester University in October and December, 1949.) [^2] Of the December edition, a participant wrote: “I wish you had been with us a few days ago we had an amusing evening discussion with Thuring [sic], Williams, Max Newman, Polyani [sic], Jefferson, J Z Young & myself $\ldots$ An electronic analyser and a digital computer (universal type) might have sorted the arguments out a bit.” Christmas postcard from Jules Y. Bogue to Warren McCulloch, c. December, 1949. American Philosophical Society, W. S. McCulloch Papers, Mss.B.M139_005. Thanks to Jonathan Swinton for this archival finding. The purpose of the proposal was to change the common meaning of the word “machine” (e.g., a steam engine, a bulldozer) in light of the new mathematical science of “universal” digital computing. The imitation game would allow for a grounded discussion of “machine” and “thinking,” seeking to expand the meaning of “thinking” and detach it from the human species, much as the meaning of “universe” was once detached from the Earth, but also as a critique of anthropocentrism.
@@
In 1950, one of the OED definitions of “machine” was: [^3] New English Dictionary. Oxford, Vol. VI, Part II, M-N, p. 7. “a combination of parts moving mechanically as contrasted with a being having life, consciousness and will $\ldots$ Hence applied to a person who acts merely from habit or obedience to a rule, without intelligence, or to one whose actions have the undeviating precision and uniformity of a machine.” Thus, by definition, common sense did not allow the meanings of “machine” and “thinking” to overlap. Despite Turing’s emphasis in his opening paragraph that he did not intend to discuss how these words were “commonly used” (p. 433), the hostility to his proposal can be seen from one of the first reactions, from a participant in the 1949 Manchester seminars, who quoted the above OED definition to appeal to common sense [mays1952, p. 149].
@@
The new question, which Turing considered to have a “more accurate form” [turing1950, p. 442], would be based on a vivid image, his “criterion for ‘thinking’ ” (p. 436), which he called interchangeably the “imitation game” and his “test.” [^4] Turing referred to his “test” four times — in pp. 446–447, 454. He also referred to it as an “experiment” — once on p. 436, twice on p. 455, and twice again on p. 457. The new question was whether a machine playing A, the deceiver, could imitate a woman, a man, a human being, or a different machine playing B, the assistant, in a remotely played conversation game to pass as B in the eyes of an average human interrogator playing C, the judge.
[…]
{mdframed}
[backgroundcolor=blue!9] BOX 1
The various questions and conditions of Turing’s test
@@
$Q^{\prime}$: “We now ask the question, ‘What will happen when a machine takes the part of A in this game?’ Will the interrogator decide wrongly as often when the game is played like this [machine-woman] as he does when the game is played between a man and a woman? These questions replace our original, ‘Can machines think?’ ” (pp. 433-434)
$Q^{\prime\prime}$: “There are already a number of digital computers in working order, and it may be asked, ‘Why not try the experiment straight away? It would be easy to satisfy the conditions of the game. A number of interrogators could be used, and statistics compiled to show how often the right identification was given.’ The short answer is that we are not asking whether all digital computers would do well in the game nor whether the computers at present available would do well, but whether there are imaginable computers which would do well.” (p. 436)
$Q^{\prime\prime\prime}$: “It was suggested tentatively that the question [$Q$] should be replaced by [$Q^{\prime\prime}$] $\ldots$ But in view of the universality property we see that either of these questions is equivalent to this, ‘Let us fix our attention on one particular digital computer C. Is it true that by modifying this computer to have an adequate storage, suitably increasing its speed of action, and providing it with an appropriate programme, C can be made to play satisfactorily the part of A in the imitation game, the part of B being taken by a man?’ ” (p. 442)
$Q^{\prime\prime\prime\prime}$: “I believe that in about fifty years’ time it will be possible to programme computers, with a storage capacity of about $10^{9}$, to make them play the imitation game so well that an average interrogator will not have more than 70 per cent. chance of making the right identification after five minutes of questioning.” (p. 442)
[…]
pink Version $Q^{\prime\prime\prime\prime}$ of the test appears at the beginning of §6 of the 1950 paper. As we will see shortly, this is the version that Turing most directly associates with the idea of a future experiment. As the least underspecified version of the test (cf. Box 1), it has been the one picked up by promoters of practical “Turing” tests. In that passage, Turing expresses his belief that in “about fifty years” an “average interrogator” would miss the identification in at least 30% of the test sessions.
Two sentences later, Turing states a second belief:
I believe that at the end of the century the use of words and general educated opinion will have altered so much that one will be able to speak of machines thinking without expecting to be contradicted. [turing1950, p. 442]
Once again we are brought to a crucial moment at the end of the century. But this second stated belief neatly reformulates the first, to which it is almost juxtaposed, and seems to reveal in common language what is meant by the rhetoric of achieving 30% of misidentification in the imitation game: it expresses in rough round numbers the experience of living in the culture he envisions, where “one will be able to speak of machines thinking without expecting to be contradicted.” Could such a cultural shift come with the future of digital computing? How remote was it? These discursive questions are arguably the real questions he addresses. What about running his test?
Having gone through nine objections to machine intelligence discussed on the basis of the imitation game, we come to the crux: Turing’s reference to “experiment” at the beginning of his §7, “Learning Machines.” This passage, which comes after his revisiting of Lady Lovelace’s Objection, has received little attention:
These last two paragraphs do not claim to be convincing arguments. They should rather be described as ‘recitations tending to produce belief.’
[Continues]
The only really satisfactory support that can be given for the view expressed at the beginning of §6 [his two stated beliefs] will be that provided by waiting for the end of the century and then doing the experiment described.
[Concludes]
But what can we say in the meantime? What steps should be taken now if the experiment is to be successful? [turing1950, p. 455]
This passage, in three sentences followed by two rhetorical questions, sums up the sophistication of Turing’s rhetoric. The suggestion to run the experiment comes juxtaposed with his indication that he is actively engaged in propaganda (“recitations tending to produce belief”). Again, he pushes “the experiment” to “the end of the century,” but now lands in the present, in a call to arms for research into “learning machines” so that the experiment — an iconic representation of the change he expects to see in talk of “machines thinking” — could be successful.
The contrast between what he proposes for the future and for the present is revealing. If the research is done as he suggests, by the time “the experiment” is to be conducted, he expects the machines to be so advanced that talk of “machines thinking” will be commonplace. In his major 1948 “Intelligent Machinery” report, [^6] Archives Centre, King’s College, Cambridge, AMT/C/11. the rhetoric of a crucial experiment does not appear. It was rather “the actual production of the machines” that “would probably have some effect” in convincing critics and opponents, because “the idea of ‘intelligence’ is itself emotional rather than mathematical” (p. 3). But in the crucial year of 1949, as we will see later, Turing faced the strongest wave of opposition from contemporaries, leading to his test.
{mdframed}
[backgroundcolor=yellow!20] BOX 2
@@
Yours sincerely, A. M. Turing777 Turing to B. H. Worsley, June 11, 1951, typeset; emphasis added. Unpublished writings of Alan Turing, copyright The Provost and Scholars of King’s College Cambridge 2023. B.H. Worsley Collection, Archives Center, National Museum of American History, Smithsonian Institution. Quoted with permission. Thanks to Mark Priestley for this archival finding.
IMITATION: FROM 1936 TO 1950
[…]
In his letter to Worsley, Turing seems to be more interested in the relations between “the motions” of Turing machines and infinite computers, whose behavior can be non-periodic. Perhaps he thought of the living human brain as an infinite computer, in the sense that it has a continuous interface with its environment, which constantly intervenes and changes its logical structure.101010“Intelligent Machinery” (op. cit.). Now, the imitation game puts into empirical form the relation between digital computers, whose behavior is ultimately periodic, and the behavior of the human players. Can the behavior of their brains be approximated by a digital computer? Turing pursued this question. For his May 1951 broadcast, he wrote: “the view which I hold myself, that it is not altogether unreasonable to describe digital computers as brains $\ldots$ If it is accepted that real brains, as found in animals, and in particular in men, are a sort of machine it will follow that our digital computer, suitably programmed, will behave like a brain.”111111“Can digital computers think?,” broadcast on BBC Third Programme, 15 May 1951. Archives Centre, King’s College, Cambridge, AMT/B/5.
{mdframed}
[backgroundcolor=pink!45] BOX 3
“…any sharp line between what machine and brain can do will fail”
@@
[…]
Yours sincerely, A. M. Turing131313 Turing to B. H. Worsley, circa June, 1951, Turing’s emphasis. Credit for this source is exactly the same as for that of Box 2. Quoted with permission.
Even if the human brain can only be compared to an infinite computer, could it not be simulated by a digital computer equipped with a sufficiently large memory? An excerpt of another newly discovered Turing letter to Worsley from mid-1951 can give more contour and provide further insight into Turing’s views (see Box 3). A highlight in this excerpt is Turing’s view that to the extent that the behavior of a neuron can be described as a stochastic process, it would be possible to “calculate random samples” of the mechanism that embodies the brain and then imitate it. An effective imitation of the brain by a machine would require knowledge of the anatomy and physiology of the brain to inspire an appropriate program, as well as much more storage and speed than was available to the Ferranti Mark I at the time (see Fig. 2). Another important element in the excerpt is Turing’s point that, even if a thinking machine is possible, the relation he has in mind is not one of identity but one of analogy: “It’ll just be another species of the thinking genus.”
Figure 2: Console of Ferranti Mark I and a group with Turing’s secretary at the Computing Machine Laboratory, Sylvia Robinson (née Wagstaff), pretending to play chess with the machine, c. 1955. Courtesy of The University of Manchester.
An original answer to the question of why design a test based on imitation, which can be seen as encouraging deception, is that imitation was actually Turing’s fundamental principle of the new science of universal digital computing. He conceived his 1950 paper partly in continuity with his 1936 paper. Both were based on his core concepts of machine and imitation, i.e., what it takes for a machine to imitate another machine. A key difference that breaks the continuity is that, by 1950, he had generalized the logical architecture of his universal machine. It would not only follow instructions, but would also be able to acquire new cognitive skills by learning. Using Turing’s 1948 language,141414“Intelligent Machinery” (op. cit.). universality can be achieved by starting with an “organized” machine (1936), or with an “unorganized” machine (1948/1950). Whereas in 1936 the machine would be given an a priori, well-defined and fixed table of instructions for each task, in 1950 it would also be able to perform a new task by changing its logical structure as a result of learning by experience, much as the brain does “by changing its neuron circuits by the growth of axons and dendrites.” [^15] Turing to Ross Ashby, circa November 19, 1946. British Library, Collection “W. Ross Ashby: correspondence of W. Ross Ashby,” Add MS 89153/26.
The use of psychological tricks in practical “Turing” tests has little to do with Turing’s 1950 proposal. In 1951, he warned: “It would be quite easy to arrange the [machine’s] experiences in such a way that they automatically caused the structure of the machine to build up into a previously intended form, and this would obviously be a gross form of cheating, almost on a par with having a man inside the machine.”161616“Intelligent machinery, a heretical theory,” a lecture given to “51 Society” at Manchester, c. 1951. Archives Centre, King’s College, Cambridge, AMT/B/4. The “human fallibility” that Turing encouraged the machine to show was meant as a by-product of learning by experience [turing1950, p. 459]: “Another important result of preparing our machine for its part in the imitation game by a process of teaching and learning is that ‘human fallibility’ is likely to be omitted [from the teaching] in a rather natural way, i.e., [learned] without special ‘coaching’.” That is, for a machine to be a valid player of Turing’s test, it cannot be specially prepared for it. This means that we have never seen a practical Turing test.
THE METHOD OF THOUGHT EXPERIMENTS
The various rhetorical questions Turing posed, $Q^{\prime}\ldots Q^{\prime\prime\prime\prime}$, to replace the original question, $Q$, can be generalized as follows [goncalves2023experiment]:
@@
It has been largely unnoticed that the various questions instantiating $Q^{\star}$ can be viewed as following a case-control methodology, applied in two stages.
[…]
yellow At the more obvious intra-game level, A plays the case, and B plays the control. However, at the inter-game level, two variants set the case (machine-woman and the machine-man) and the other two set the control (man-woman and the machine-machine). While the first two are open, creating suspense around the test, the latter two are resolved as follows. It is known that a man (A) can imitate gender stereotypes associated with a woman (B) to deceive an interrogator (C) despite their physical differences, as assumed in parlor games and as Turing illustrates on p. 434: “My hair is shingled…” Further, regarding the machine-machine variant, it is also known that a digital computer (A), because of its universality property [turing1950, §§4, 5], can imitate any discrete-state machine (B) and even a continuous one (p. 451), despite their physical differences.
We can now explore how Turing’s presentation of his test fits Mach’s description of “the basic method of thought experiments,” which is variation, continuously if possible. Mach is the author of perhaps the most classic text on thought experiments in the modern scientific tradition [mach1897], in which he developed observations and insights based on examples from the history of modern physics, mathematics, and common sense experience. He wrote: “By varying the conditions (continuously if possible), the scope of ideas (expectations) tied to them is extended: by modifying and specializing the conditions we modify and specialize the ideas, making them more determinate, and the two processes alternate” (p. 139). Mach illustrated his point with the process of discovery of universal gravitation:
A stone falls to the ground. Increase the stone’s distance from the earth, and it would go against the grain to expect that this continuous increase would lead to some discontinuity. Even at lunar distance the stone will not suddenly lose its tendency to fall. Moreover, big stones fall like small ones: the moon tends to fall to the earth. Our ideas would lose the requisite determination if one body were attracted to the other but not the reverse, thus the attraction is mutual and remains so with unequal bodies, for the cases merge into one another continuously $\ldots$ discontinuities are quite conceivable, but it is highly improbable that their existence would not have betrayed itself by some experience. [mach1897, pp. 138-139]
@@
Turing’s imitation game extended the scope of ideas and expectations established earlier in his 1936 paper, moving from machine-machine and restricted human-machine imitation in 1936171717“We may compare a man in the process of computing a real number to a machine which is only capable of a finite number of conditions” [turing1936, p. 231]. to more general human-machine imitation in 1950.